One-third of the sum of two consecutive integers is 9 less than the smaller integer. Find the integers.
Answer by Nate(3500)

first integer = i
second integer = i + 1
(1/3)(i + i + 1) = i - 9
2i + 1 = 3i - 27
28 = i
first integer = 28
second integer = 29